Calculate the velocity magnitude and angle of the PROJECTILE listed below

If it's a projectile and I get a negative angle, is that ok?
8.1 vy
Answer #1

According to the question we need to find the magnitude and direction ( Angle ) .

V(mag ) = sqrt ( v(x)^2 + v(y) ^2 ) = sqrt ( 8.1 ^2 + ( -7.2 ) ^2 ) = 10.837435 m/s ======Part a)

Angle = tan ^-1 ( v(y) / v (x) ) = tan ^-1 ( -7.2 / 8.1 ) = -41.63353 deg ===============Part b)